OpenStudy (anonymous):

r+s=12 4r-6s=12 I need to solve the system by substitution and check my answer, please help :(((

OpenStudy (mertsj):

1. Solve the first equation for r 2. Replace r in the second equation with the expression you got in Step 1 3. Solve the resulting equation for s 4. Replace s with that number in an equation that has both s and r. 5. Solve for r

OpenStudy (anonymous):

can you do it....

OpenStudy (anonymous):

like the first one so i can see

OpenStudy (mertsj):

1. r=12-s 2. 4(12-s)-6s=12 48-4s-6s=12 -10s=-36 s=3.6 3. r=12-3.6 4. r=8.4

OpenStudy (anonymous):

x+y=12 x-y=2 solve by elimination

OpenStudy (mertsj):

Ok. I did the first one so you could see. Now it's your turn.

OpenStudy (anonymous):

but this one is elimanation

OpenStudy (mertsj):

Add the two equations.